OpenStudy (anonymous):

Simplify the Rational Expression:

OpenStudy (anonymous):

\[\frac{ 3x+2 }{ 15x2+16x+4 }\]

mathslover (mathslover):

Ok so first of all factorize the denominator

mathslover (mathslover):

can you?

OpenStudy (anonymous):

Yes

mathslover (mathslover):

tell me what you get @GenGen96 ?

OpenStudy (anonymous):

(x+6)(x+10)

mathslover (mathslover):

do you mean to say that; \[\large{(x+6)(x+10)=15x^2+16x+4}\] ?

mathslover (mathslover):

check it again @GenGen96

OpenStudy (anonymous):

\[(5x+2)(3x+2)\]

mathslover (mathslover):

good now.. we have \[\large{\frac{3x+2}{(5x+2)(3x+2)}}\] can you cancel out like term?

OpenStudy (anonymous):

So would it just be \[\frac{ 1 }{ (5x+2) }\]

mathslover (mathslover):

right

mathslover (mathslover):

good work

OpenStudy (anonymous):

Can you help me find all the numbers that much be excluded from the domain?

mathslover (mathslover):

Post that as a diff. question that will be answered there I am sure